# Detecting Subtle Differences between Human and Model Languages Using Spectrum of Relative Likelihood

### v1

@@ Abstract
Yang Xu1,*, Yu Wang2, Hao An1, Zhichen Liu1, Yongyuan Li1, 1Department of Computer Science and Engineering, Southern University of Science and Technology 2Bielefeld University, *Correspondence: xuyang@sustech.edu.cn
@@ 1 Introduction
One of the recent endeavors in natural language generation is to develop effective methods of detecting model-generated texts from real human texts. The current most effective methods for this task utilize the likelihood information in text data. More specifically, most existing methods rely on the observation that the absolute likelihood values of texts naturally distribute differently, depending on their sources. Examples include the early work like GLTR (Gehrmann et al., 2019) and more recent ones like DetectGPT (Mitchell et al., 2023) and Fast-DetectGPT (Bao et al., 2024).
However, existing methods have the following limitations: First, most work estimates likelihood as a static property of language, but overlooks the fact that human language processing is a dynamic process during which the likelihood of information under processing is bound to certain linguistic and cognitive constraints. For example, the trade-off between processing effort and likelihood of words (Levy, 2008; Smith and Levy, 2013), limited attention focus, activation decay Lewis et al. (2006) and so on. Second, merely using the absolute likelihood values to distinguish generated and human texts is a tricky “cat and mouse game” – as models’ capability of mimicking human language keeps growing, their productions would eventually become hardly distinguishable. Third, current methods are not computationally economical, because most of them need to run at least one time of inference on text with a fairly large language model.
We propose a human-model text detection approach that addresses the aforementioned issues and achieves better or competitive performance with existing methods. Our approach extracts features from the spectrum view of relative likelihood scores of texts, to capture the dynamic changes of likelihood in language. These features are used to design two types of classifiers, a supervised learning-based one and a heuristic-based zero-shot one, both of which reach impressive performances. The core idea is to obtain the spectrum of likelihood using the Fourier transform, which summarizes the complex patterns of likelihood change in time domain into a much more compact view that magnifies the subtle differences between different texts. It has a theoretical basis in psycholinguistic studies on surprisal (likelihood) distribution in natural language. Further, our method is still effective when likelihood scores are estimated by naïve n-gram models, which places much less computational cost. We name our approach FourierGPT, inspired by existing methods like DetectGPT and Fast-DetectGPT.
@@ 2.1 Likelihood-based zero-shot text detection
Two other likelihood-based methods are also compared within this study: normalized log-rank perturbation (NPR) (Su et al., 2023) and divergence between multiple completions of a truncated passage (DNA-GPT) (Yang et al., 2023a). Both rely on estimation of absolute likelihood to some extent.
@@ 2.2 Surprisal and likelihood of language
Understanding the human mind’s preference and tendency in handling surprisal/likelihood leads to new ideas for natural language generation techniques. For example, some recent endeavors build on top of the assumption that model-generated language appears more natural and human-like if it is generated through a decoding algorithm that follows the UID theory, such as the beam search algorithm as evidenced in Meister et al. (2020); or it falls under the so-called stable entropy zone (Arora et al., 2023); Meister et al. (2023) propose locally typical sampling, which enforces the uniform distribution of likelihood during the generation process, and results in generated texts that are more aligned with human texts.
@@ 2.3 Evaluation of natural language generation with likelihood
Some recent evaluation metrics compare model text with human text in high dimensional space, such as MAUVE (Pillutla et al., 2021). While this type of method does not directly use likelihood information, interesting correlations have been found between likelihood-based metrics. For example, Yang et al. (2023b) proposes a novel evaluation metric called Fourier analysis of cross-entropy (FACE), which converts the cross-entropy scores (i.e., likelihood) to spectrum representations and then measures the distances in frequency-domain. The resulted distance scores can reflect generation qualities that are co-examined with other metrics, such as MAUVE, and align well with human judgements. This work indicates that with proper transformation on simple likelihood scores, rich insights about language use are viable. In fact, the text detection method proposed in this study is directly inspired by Yang et al.’s work (2023b).
@@ 3.1 Estimation and normalization of likelihood scores
We estimate the likelihood scores of text data with pretrained language models of various scales: Mistral-7B (Jiang et al., 2023), GPT-2 families (Radford et al., 2019), and a bigram language model trained from scratch. We use the implementation of bigram model from Heafield (2011) and train it on a subset of C4 dataset (Raffel et al., 2019). Raw likelihood scores are estimated by running a forward pass on the input text. Taking an input of $N$ tokens, $t_{1},\dots,t_{N}$, the likelihood score of $i$th token $s_{i}$ is computed by $s_{i}=-\log P(t_{i}|t_{1},\dots,t_{i-1})$, i.e., the negative log probability returned by the language model, which we call the estimator model.
Then we normalize the raw likelihood scores $s_{1},\dots,s_{N}$ within each sequence, obtaining the $z$-scored likelihood $\tilde{s}_{1},\dots,\tilde{s}_{N}$, in which $\tilde{s}_{i}=\frac{s_{i}-\mu}{\sigma}$, $\mu=\sum s_{i}/N$ and $\sigma=\sum(s_{i}-\mu)^{2}/(N-1)$. Here we would like to stress that this seemingly trivial normalization step is actually critical for verifying the hypothesis of this study: the raw likelihood $s_{i}$’s value depends on the choice of estimator – larger models usually result in smaller values (similar to perplexity); but the $z$-scored $\tilde{s}_{i}$ characterizes the relative level of likelihood within the range $(0,1)$, which is less dependent on the estimator. As expected, the distribution of raw likelihood is highly skewed, while the $z$-scored likelihood is closer to normal distribution and gives better classification results in the following steps.
@@ 3.2 Fourier transform
The range of $\omega_{k}$ is $[0,\pi]$, and its interpretation is not trivial. Based on an intuitive interpretation provided by Yang et al. (2023b), we can roughly tell that the likelihood score $\tilde{s}_{i}$ at the level of $\lVert X(\omega_{k})\rVert$ tends to occur every $1/\omega_{k}$ tokens in the text data. Interestingly, we find the way $\lVert X(\omega_{k})\rVert$ distributes along $\omega_{k}$ provides unique information to distinguish human from model. To develop solid explanations of what the spectrum of likelihood means is important, yet a different topic. We primarily focus on harnessing the spectrum information for the detection task, and try to do gain some interpretive insights at our best in Section 5.
@@ 4.1 Datasets
Table 2: Accuracy of our best supervised classifiers compared to other likelihood-based zero-shot methods reported in (Bao et al., 2024) on selected task subsets. Best scores are in bold, and $\dagger$ indicates second best.
@@ 5.1 Answers starting with “yes/no”
We use a simple ablation experiment on data to examine whether this subtle difference is reflected in the spectrum of likelihood. We remove the “Yes”/“No” at the beginning of the answer, re-computing the likelihood scores, and re-do the Fourier transform. Consequently, the spectrum of the model morphs in shape towards the direction of human (Figure 3 (left)): the altered GPT-4 data’s low-frequency components drop, and the high ends rise, both towards the direction of human. To showcase the advantage of spectrum view, we also plot the $z$-scored likelihood against token position (Figure 3 (right)), which shows that removing “Yes”/“No” makes the likelihood curve flatter (thus, closer to human), but this change is not as easy to describe as the spectrum. In sum, the subtle differences between human and model languages, like the “Yes”/“No” use discussed here, can be reflected in likelihood space, and the spectrum view can capture this difference conveniently.
Figure 3: The changes of likelihood spectrum (left) and likelihood-position plot (right) after removing the “Yes”/“No” in answer from PubMed data (with GPT-4 only). Curves are fit with GAM. Shaded areas are 95% confidence intervals from bootstrap.
@@ 5.2 Text lengths effect
Figure 4: Text lengths affect likelihood spectrum and pairwise classifier performance. Each plot corresponds to lengths of text, $n=50,100,150$, compared to “Full”.
@@ 5.3 Part-of-speech masking
As the last part of discussion, we test the role played by words of different part-of-speech (POS) tags in affecting the likelihood spectrum. First, we mask three POS tags in text: ‘NOUN’, ‘VERB’, and ‘ADJ’, individually; and the union of the three, ‘NOUN+VERB+ADJ’ (NVA). Then, the masked tokens’ likelihood scores are replaced with the average score, thus eliminating the contribution from that specific POS tag. Masking is done use the spaCy (Honnibal et al., 2020) POS tagger. The likelihood spectrum results after ‘VERB’ being masked is shown in Figure 5.
(a) Xsum
(b) PubMed
(c) Writing
Figure 5: Likelihood spectrum before and after masking on ‘VERB’ tag with GPT-2 as estimator.
We find an interesting phenomenon that, after applying the POS mask, the change of likelihood spectrum for human text is relatively small, while for the model text (GPT-4), the change is much bigger (see (a) and (c) in Figure 5). Such difference is not limited to ‘VERB’, but also observed for the other POS tags (shown from Figure 6 to Figure 8 in Appendix). The difference will be more salient if switching the estimator from GPT-2 to bigram (shown in Figure 9 to Figure 12 in Appendix). We back up the observed difference by calculating the spectral overlap, the metric proposed by Yang et al. (2023b) to measure the similarity between two spectra: The original spectrum and the one after POS masking, for both human and model texts. It turns out that indeed model text has smaller spectral overlap compared to human text, which indicates that the model spectrum is less stable against POS masking. (Details reported in Table 8 in Appendix)
This finding is similar to the finding about probability curvature in DetectGPT Mitchell et al. (2023): human text is less likely to reach the local maxima of likelihood than model-generated text. In our case: there is more randomness in real human text and as a result, it is more stable against perturbation of likelihood in time-domain, such as the POS masking. Therefore, spectrum stability is a potential statistic for zero-shot detection.
@@ 7 Limitations
The limitations of the current study are: First, the pair-wise classifier requires the two texts being classified must be generated from the same prompt. It is yet to be verified whether the classifier’s performance will retain if the source prompts are different. Second, the supervised classifier is not a strict zero-shot detector, and its performance still has space for improvement. We argue that it would be less of an issue if the classifier learns the general features instead of idiosyncrasy in certain data. This, however, requires further investigation to whether likelihood spectrum is such a general feature. Third, the datasets examined are relatively small. It is worth exploration on larger datasets (especially short text corpus, such as QA) to further confirm the effectiveness of the method.
@@ Appendix A Hyperparameters for Supervised Classifiers
Table 8: Attention mask effect on Spectral Overlap. H denotes human. M denotes model.Bolded Number is the minority which shows that likelihood spectrum from human text changes more than the equivalent from model generated text after attention mask.All the other number shows the opposite situation

### v2

@@ Abstract
Yang Xu1,*, Yu Wang2, Hao An1, Zhichen Liu1, Yongyuan Li1 1Department of Computer Science and Engineering, Southern University of Science and Technology 2Digital Linguistics Lab, Department of Linguistics, Bielefeld University *Correspondence: xuyang@sustech.edu.cn
@@ 1 Introduction
A recent focus in natural language generation is to develop effective methods of detecting model-generated texts from real human texts. This endeavor gives human users an opportunity to know whether a given text, e.g., an explanation, is either generated by language model or written by human, allowing them to decide how much trust to place in the explanation. The current most effective methods for this task utilize the likelihood information in text data (e.g.,see Verma et al., 2024; Venkatraman et al., 2024). More specifically, most existing methods rely on the observation that the absolute likelihood values of texts naturally distribute differently, depending on their sources. Examples include the early work like GLTR (Gehrmann et al., 2019) and more recent ones like DetectGPT (Mitchell et al., 2023) and Fast-DetectGPT (Bao et al., 2024).
However, existing methods have the following limitations: First, most work estimates likelihood as a static property of language, but overlooks the fact that human language processing is a dynamic process during which the likelihood of information under processing is bound to certain linguistic and cognitive constraints. For example, the trade-off between processing effort and likelihood of words (Levy, 2008; Smith and Levy, 2013), limited and selective context access (Gibson, 1998, 2000), activation decay Lewis et al. (2006) and so on. Second, merely using the absolute likelihood values to distinguish generated and human texts is a tricky “cat and mouse game” – as models’ capability of mimicking human language keeps growing, their productions would eventually become hardly distinguishable (Brown et al., 2020). Third, current methods are not computationally economical, because most of them need to run at least one time of inference on text with a fairly large language model.
We propose a human-model text detection approach that addresses the aforementioned issues and achieves better or comparable performance with existing methods. Our approach extracts features from the spectrum view of relative likelihood scores of texts, to capture the dynamic changes of likelihood in language. These features are used to design two types of classifiers, a supervised learning-based one and a heuristic-based zero-shot one, both of which reach impressive performances. The core idea is to obtain the spectrum of likelihood using the Fourier transform, which summarizes the complex patterns of likelihood change in time domain into a much more compact view that magnifies the subtle differences between different texts. It has a theoretical basis in psycholinguistic studies on surprisal (likelihood) distribution in natural language. Further, our method is still effective when likelihood scores are estimated by naïve n-gram models, which places much less computational cost. We name our approach FourierGPT, inspired by existing methods like DetectGPT and Fast-DetectGPT.
@@ 2.1 Likelihood-based zero-shot text detection
Two other likelihood-based methods are also compared within this study: normalized log-rank perturbation (NPR) (Su et al., 2023) and divergence between multiple completions of a truncated passage (DNA-GPT) (Yang et al., 2024). Both rely on estimation of absolute likelihood to some extent.
@@ 2.2 Surprisal and likelihood of language
Understanding the human mind’s preference and tendency in handling surprisal/likelihood leads to new ideas for natural language generation techniques. For example, some recent endeavors build on top of the assumption that model-generated language appears more natural and human-like if it is generated through a decoding algorithm that follows the UID theory, such as the beam search algorithm as evidenced in Meister et al. (2020); or it falls under the so-called stable entropy zone (Arora et al., 2023); Meister et al. (2023) proposes locally typical sampling, which enforces the uniform distribution of likelihood during the generation process, and results in generated texts that are more aligned with human texts.
@@ 2.3 Evaluation of natural language generation with likelihood
Some recent evaluation metrics compare model text with human text in high dimensional space, such as MAUVE (Pillutla et al., 2021). While this type of method does not directly use likelihood information, interesting correlations have been found between likelihood-based metrics. For example, Yang et al. (2023) proposes a novel evaluation metric called Fourier analysis of cross-entropy (FACE), which converts the cross-entropy scores (i.e., likelihood) to spectrum representations and then measures the distances in frequency-domain. The resulted distance scores can reflect generation qualities that are co-examined with other metrics, such as MAUVE, and align well with human judgements. This work indicates that with proper transformation on simple likelihood scores, rich insights about language use are viable. In fact, the text detection method proposed in this study is directly inspired by Yang et al.’s work (2023).
@@ 3.1 Estimation and normalization of likelihood scores
We estimate the likelihood scores of text data with pretrained language models of various scales: Mistral-7B (Jiang et al., 2023), GPT-2 families (Radford et al., 2019), and a bigram language model trained from scratch. We use the implementation of bigram model from Heafield (2011) and train it on a subset of C4 dataset (Raffel et al., 2020). Raw likelihood scores are estimated by running a forward pass on the input text. Taking an input of $N$ tokens, $t_{1},\dots,t_{N}$, the likelihood score of $i$th token $s_{i}$ is computed by $s_{i}=-\log P(t_{i}|t_{1},\dots,t_{i-1})$, i.e., the negative log probability returned by the language model, which we call the estimator model.
Then we normalize the raw likelihood scores $s_{1},\dots,s_{N}$ within each sequence, obtaining the $z$-scored likelihood $\tilde{s}_{1},\dots,\tilde{s}_{N}$, in which $\tilde{s}_{i}=\frac{s_{i}-\mu}{\sigma}$, $\mu=\sum s_{i}/N$ and $\sigma=\sum(s_{i}-\mu)^{2}/(N-1)$. Here we would like to stress that this seemingly trivial normalization step is actually critical for verifying the hypothesis of this study: the raw likelihood $s_{i}$’s value depends on the choice of estimator – larger models usually result in smaller values (similar to perplexity); but the $z$-scored $\tilde{s}_{i}$ characterizes the relative level of likelihood within the range (,)($\infty$,$\infty$)( - ∞ , ∞ ), which is less dependent on the estimator. As expected, the distribution of raw likelihood is highly skewed, while the $z$-scored likelihood is closer to normal distribution and gives better classification results in the following steps.
@@ 3.2 Fourier transform
The range of $\omega_{k}$ is $[0,\pi]$, and its interpretation is not trivial. Based on an intuitive interpretation provided by Yang et al. (2023), we can roughly tell that the likelihood score $\tilde{s}_{i}$ at the level of $\lVert X(\omega_{k})\rVert$ tends to occur every $1/\omega_{k}$ tokens in the text data. Interestingly, we find the way $\lVert X(\omega_{k})\rVert$ distributes along $\omega_{k}$ provides unique information to distinguish human from model. To develop solid explanations of what the spectrum of likelihood means is important, yet a different topic. We primarily focus on harnessing the spectrum information for the detection task, and try to do gain some interpretive insights at our best in Section 5.
@@ 4.1 Datasets
Table 2: Accuracy of our best supervised classifiers compared to other likelihood-based zero-shot methods reported in (Bao et al., 2024) on selected two task subsets. Best scores are in bold, and $\dagger$ indicates second best.
@@ 5.1 Answers starting with “yes/no”
We use a simple ablation experiment on data to examine whether this subtle difference is reflected in the spectrum of likelihood. We remove the “Yes”/“No” at the beginning of the answer, re-computing the likelihood scores, and re-do the Fourier transform. Consequently, the spectrum of the model morphs in shape towards the direction of human (Figure 3(a) (left)): the altered GPT-4 data’s low-frequency components drop, and the high ends rise, both towards the direction of human. GPT-3.5 and GPT-3 have much fewer data points containing “Yes”/“No”, but a similar trend exists in the high end (see Figure 3(b)). To showcase the advantage of spectrum view, we also plot the $z$-scored likelihood against token position (Figure 3(a) (right)), which shows that removing “Yes”/“No” makes the likelihood curve flatter (thus, closer to human), but this change is not as easy to describe as the spectrum. In sum, the subtle differences between human and model languages, like the “Yes”/“No” use discussed here, can be reflected in likelihood space, and the spectrum view can capture this difference conveniently.
(a) Estimated by GPT-4.
(b) Estimated by all three models.
Figure 3: (a) The changes of likelihood spectrum ((a) left) and likelihood-position plot ((a) right) after removing the “Yes”/“No” in answer from PubMed data (estimated by GPT-4 only). (b) The changes of spectrum estimated by all three models. Curves are fit with GAM. Shaded areas are 95% confidence intervals from bootstrap.
@@ 5.2 Text lengths effect
Figure 4: Text lengths affect likelihood spectrum and pairwise classifier performance. Each plot corresponds to lengths of text, $n=50,100,150$, compared to “Full”. Curves are fit with GAM. Shaded areas are 95% confidence intervals from bootstrap.
@@ 5.3 Part-of-speech masking
As the last part of discussion, we investigate the role played by words of different part-of-speech (POS) tags in affecting the likelihood spectrum. Many previous studies have revealed that adding POS information in either the encoder or decoder side of (multimodal) language models can enhance language understanding and generation tasks (e.g., see Wray et al., 2019; Deshpande et al., 2019). Yang et al. (2022) further demonstrated that by informing the sampling method of different POS patterns, the generated texts show more diversity without sacrificing quality. The above findings suggest that, POS patterns as linguistic knowledge can significantly impact the NLG process.
Based on this, we use masking to break some selected POS patterns during the generation process, which will restrain the estimated likelihood from encoding complete information of POS, and then see how this will affect the likelihood spectrum. First, we mask three POS tags in text: ‘NOUN’, ‘VERB’, and ‘ADJ’, individually; and the union of the three, ‘NOUN+VERB+ADJ’ (thus, ‘NVA’). Then, the masked tokens’ likelihood scores are replaced with the average score, thus eliminating the contribution from that specific POS tag. Masking is done use the spaCy (Honnibal et al., 2020) POS tagger. The likelihood spectrum results after ‘NVA’ being masked is shown in Figure 5.
(a) PubMed
(b) Writing
(c) Xsum
Figure 5: Likelihood spectrum before and after masking out the ‘NOUN+VERB+ADJ’ POS tags. It can be seen that GPT-4 texts show greater variation before and after the mask, while human text show smaller change. Likelihood is estimated with a bigram model. Curves are fit with GAM. Shaded areas are 95% confidence intervals from bootstrap.
We find an interesting phenomenon: after applying the POS mask, the change of likelihood spectrum for human text is relatively small, while for the model text (GPT-4), the change is much bigger (see Figure 5). Such difference is not limited to the collection of ‘NVA’ tags alone, but also observed for the ‘NOUN’, ‘VERB’, and ‘ADJ’ POS tags individually (shown from Figure 6 to Figure 8 in Appendix). This difference is most salient if bigram is used as the estimator [^2] Technically, it is impossible to apply attention mask in bigram model. Our solution is: as an example, if a verb is encountered, we replace the current likelihood value with a random value between the minimum and maximum, both of which are computed in the same sentence where the verb is located. This operation metaphorically mimics the attention masking. as compared to GPT-2. We back up the observation by calculating the spectral overlap between the original spectrum and the one after POS masking, for both human and model texts. It is a common metric for measuring spectrum similarity, and it was first used by Yang et al. (2023) as a means for NLG evaluation. It turns out that indeed the generated text has smaller before-after spectral overlap, as compared to human text, which indicates that the model spectrum is less stable against POS masking (Details reported in Table 8 in Appendix).
This finding is consistent with the main finding about probability curvature in DetectGPT (Mitchell et al., 2023): human text is less likely to reach the local maxima of likelihood than model-generated text. In our case, the results suggest that there is more randomness in real human text, which makes it more stable against perturbation in time domain, such as the POS masking.
@@ 7 Limitations
The limitations of the current study are: First, the pair-wise classifier requires that the two texts being classified must be generated from the same prompt. It is yet to be verified whether the classifier’s performance will retain if the source prompts are different. Second, the supervised classifier is not a strict zero-shot detector, and its performance still has space for improvement. We argue that it would be less of an issue if the classifier learns the general features instead of idiosyncrasy in certain data. This, however, requires further investigation to whether likelihood spectrum is such a general feature. Third, the datasets examined are relatively small. It is worth exploration on larger datasets (especially short text corpus, such as QA) to further confirm the effectiveness of the method.
Acknowledgments
We sincerely thank all the reviewers for their efforts in pointing out the weakness in the paper and their insightful advice for future improvement. Yu Wang is supported by the Deutsche Forschungsgemeinschaft (DFG): TRR 318/1 2021 – 438445824.
@@ Appendix A Hyperparameters for Supervised Classifiers
Table 8: Attention mask effect on Spectral Overlap. H denotes human. M denotes model.Bolded Number is the minority which shows that likelihood spectrum from human text changes more than the equivalent from model generated text after attention mask. All the other number shows the opposite situation.